OpenStudy (anonymous):

Prove that any complex number with modulus 1 (exept -1) can be represented in the form: (1+it)/(1-it) where t is a real

OpenStudy (experimentx):

both numerator and denominator has the same moduli (1+it)/(1-it) ... so this term has modulus 1

OpenStudy (anonymous):

\[\frac{1+it}{1-it} \frac{1+it}{1+it}=\frac{1-t^2+2ti}{1+t^2}=\frac{1-t^2}{1+t^2}+\frac{2t}{1+t^2} i\]

OpenStudy (anonymous):

and?

OpenStudy (experimentx):

or any z = a + ib such that a^2 + b^2 = 1 |1 - it| = |1 + it| = sqrt(1 + t^2) arg(1 - it) = - arg(1 + it) e^(2arg(1+it)) = sqrt(a^2 + b^2)e^(2*(1/2)arg(z))) = sqrt(a^2+b^2)e^(1/2*arg(z))/(sqrt(a^2+b^2)e^(-1/2 arg(z))) \ \[ \frac{\sqrt{a^2 + b^2}e^{{1\over 2}arg(z)}}{\sqrt{a^2 + b^2}e^{{-1\over 2}arg(z)}} = \frac{\sqrt{1 + (b/a)^2}e^{{1\over 2}arg(z)}}{\sqrt{1 + (-b/a)^2}e^{{-1\over 2}arg(z)}} \]

OpenStudy (experimentx):

still that 1/2 is troubling.

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

tx @experimentX

OpenStudy (anonymous):

Oo

OpenStudy (experimentx):

hold on ... \[ \huge a + ib = e^{i\arctan(b/a)} =\\ \huge \frac{1 + i {\sin (1/2 \arctan(b/a))\over \cos (1/2 \arctan(b/a))}}{ 1 - i {\sin (1/2 \arctan(b/a))\over \cos (1/2 \arctan(b/a))}} \]

OpenStudy (anonymous):

\[\frac {\sqrt{1+t ^{2}}e ^{\frac{1}{2}argz}}{\sqrt{1+t ^{2}}e ^{-\frac{1}{2}argz}}=e ^{\arg(z)}=z\]

OpenStudy (anonymous):

thx again

OpenStudy (experimentx):

For -1, if you use the above formula, you get -1 = 1

OpenStudy (experimentx):

no probs.